Can someone please tell me what are the following formulas for:
1.) the hypotenuse of a 45-45-90 isosceles right triangle with a leg of length x.
2.) the hypotenuse of a 30-60-90 right triangle with shortest leg having a length of x.
3.) the cross section through the center of a sphere.
4.) the volume of a prism that has an equilateral triangle with side of length x and height of length y.
I believe that it is asking for the formulas only. These are the last four i need. so please help! thank you!

Answer by stanbon(48546)

what are the following formulas for:
1.) the hypotenuse of a 45-45-90 isosceles right triangle with a
leg of length x.
Draw the figure. Label the legs "x". Label the hypotenuse "h".
h=sqrt(x^2 + x^2)
--------------------------------
2.) the hypotenuse of a 30-60-90 right triangle with shortest leg having a length of x.
Draw the figure. Label the short side as "x". Thenthe hypotenuse is h=2x
and the 3rd side is xsqrt3.
h=sqrt(x^2 + 3x^2)
---------------------------
3.) the cross section through the center of a sphere.
The cross section is a circle:
x^2+y^2=1
----------------------------
4.) the volume of a prism that has an equilateral triangle with side of length x and height of length y.
What do you mean by "has an equilateral triangle"?
Volume requires three dimensions. Where is the 3rd dimension?
----------------------------
Cheers,
Stan H.